# MS-0000-0000.00

### v1

@@
Artificial Intelligence (AI) is advancing at an unprecedented pace, with clear potential to enhance decision-making and productivity. Yet, the collaborative decision-making process between humans and AI remains underdeveloped, often falling short of its transformative possibilities. This paper explores the evolution of AI agents from passive tools to active collaborators in human-AI teams, emphasizing their ability to learn, adapt, and operate autonomously in complex environments. This paradigm shift challenges traditional team dynamics, requiring new interaction protocols, delegation strategies, and responsibility distribution frameworks. Drawing on Team Situation Awareness (SA) theory, we identify two critical gaps in current human-AI teaming research: the difficulty of aligning AI agents with human values and objectives, and the underutilization of AI’s capabilities as genuine team members. Addressing these gaps, we propose a structured research outlook centered on four key aspects of human-AI teaming: formulation, coordination, maintenance, and training. Our framework highlights the importance of shared mental models, trust-building, conflict resolution, and skill adaptation for effective teaming. Furthermore, we discuss the unique challenges posed by varying team compositions, goals, and complexities. This paper provides a foundational agenda for future research and practical design of sustainable, high-performing human-AI teams.
@@
Human-AI Teaming; Situation Awareness; Research Outlook
@@ 1 Introduction
Artificial Intelligence (AI) is advancing at an unprecedented pace, transforming industries and reshaping daily life. From healthcare and finance to education and manufacturing, AI-based systems are increasingly supporting humans across various fields (Lu and Zhang 2022, Lou and Wu 2021, Berente et al. 2021). Yet, despite AI’s potential, collaborations between humans and AI remain ad hoc, often resulting in underwhelming performance (Teaming 2022). The promise of AI lies in its ability to integrate into teams to augment human capabilities, improve decision-making, enhance productivity, and drive innovation. Realizing the benefits, however, requires a more systematic examination of the novelties associated with human-AI teaming.
@@
In traditional human-human teaming, researchers have extensively studied shared awareness, communication, and collaboration, establishing frameworks around communication frequency, shared mental models, task complexity, and training. Similarly, human-system teaming has focused on the use of communication and collaboration tools, which support team coordination but remain distinct from the tasks themselves (DeSanctis and Gallupe 1987, Jerry Fjermestad 2000, Dennis and Wixom 2002). Implicit in all team-based interactions is the assumption that team members share common values and work together toward shared goals and objectives. The perspective that technology is a tool to enhance human efforts needs a refresh with the emergence of AI agents. Recent studies have shown that AI agents are evolving rapidly, with their capabilities doubling approximately every seven months since 2019 (Kwa et al. 2025). [^1] The concept of AI as agents is not new. Russell and Norvig (2016) define AI agents as entities that perceive their environment and act rationally. This paper builds on and modernizes that foundational view to reflect the advanced capabilities of today’s AI. We explore how AI agents have evolved beyond mere rational actors to become autonomous, iterative, and intelligent team members. These agents now exhibit the ability to learn, reflect, and adapt within complex and dynamic environments. Moreover, we highlight a significant transformation: AI agents are no longer limited to serving as passive tools but have emerged as active collaborators in human-AI teams. This evolution marks a paradigm shift, as AI agents now operate in contexts demanding emotional intelligence, sustained interaction, and autonomous problem-solving—attributes that transcend the boundaries of the classical conceptualization of AI agents. In this regard, human-AI teaming introduces a new era of team science, requiring a rethinking of interaction protocols for AI agents, as well as new approaches to delegation, task execution, and the distribution of responsibilities in team processes.
@@
In this paper, we examine the emerging landscape of human-AI teaming, where AI has moved beyond simple automation to become an intelligent, agentic artifact (Dennis et al. 2023, Gao et al. 2024b). The rapid rise of large language models (LLM) such as GPT has accelerated this transformation. AI-enabled agentic artifacts are no longer simply embedded into workflows; they are now intelligent and iterative agents capable of actively taking on mission-critical roles and tasks within teams (McKinsey 2023, 2024, Alavi 2024). These systems learn from interactions, adapt to changing contexts, and make independent decisions, transforming AI into an autonomous partner in collaborative efforts. The agentic shift highlights the gap in the current information systems literature, which mainly treats AI as a tool rather than a nearly full-fledged teammate (Wang et al. 2019, Dennis et al. 2023, Lu and Zhang 2022).
@@
This paper aims to bridge these gaps by applying the well-established Team Situation Awareness (SA) theory to the context of human-AI teaming. We argue that integrating AI agents into teams not only requires designing systems that can work toward shared objectives but also requires a rethinking of how teams execute tasks alongside AI agents. By doing so, we identify significant gaps in how Team SA has been applied to AI, and we propose a structured outlook for future research to address these issues. Furthermore, we propose a structured framework for understanding human-AI teaming across four key dimensions: team formulation, coordination, social dynamics, and knowledge creation. First, effective team formulation requires shared mental models that align AI’s decision-making with human values and objectives, ensuring clarity in role specification while allowing flexibility for adaptation. Team coordination focuses on optimizing task delegation and human-AI interaction protocols, minimizing expectation misalignments and ensuring interpretable and reliable decisions. Team maintenance is crucial for trust-building, conflict resolution, and sustaining high-performance collaboration, particularly in managing AI’s limitations, biases, and accountability in decision-making. Finally, team training and evolution involve continuous learning mechanisms that enable AI to refine its capabilities while ensuring human members continue to create new knowledge. By structuring research around these four interconnected areas, we outline a roadmap for advancing human-AI teaming and addressing critical challenges such as trust, adaptability, accountability, and long-term sustainability in AI-integrated teams.
@@ 2.1 Emerging Agentic AI Systems
Information systems (IS) are beginning their transformation into agentic artifacts that “are no longer passive tools waiting to be used, are no longer always subordinate to the human agent, and can now assume responsibility for tasks with ambiguous requirements and for seeking optimal outcomes under uncertainty” (Baird and Maruping 2021). Modern AI systems, with their ability to “learn and act” independently of human instructions are showing the promise to emulate human thought processes, behavior, and decision-making (Dennis et al. 2023, Durante et al. 2024). Although these systems may remain within the realm of Narrow AI, their capabilities now extend to many tasks that, until recently, required highly trained professionals.
Generative AI systems, such as GPT, Midjourney, Claude, Llama, and Gemini, can perform tasks requiring nuanced judgment, and adaptability to novel contexts — areas where classical AI systems often falter due to “brittleness” when encountering scenarios outside their predefined rules (McKinsey 2023, Dhar 2023). Generative AI tools now assist in domains previously dominated by human cognition and expertise (Zhou and Lee 2024, Susarla et al. 2023). For example, ChatGPT can engage in meaningful conversations, provide detailed answers, and even draft complex documents by interpreting ambiguous prompts. Similarly, Midjourney can create highly detailed and original artwork from simple text descriptions, demonstrating creative problem-solving skills that go beyond rule-based automation.
Agentic systems operate in collaborative environments where they not only participate in tasks requiring creativity and coordination but also take initiative in decision-making. Rather than following static scripts, they analyze diverse inputs, generate contextually appropriate outputs, and continuously improve through iterative feedback. In doing so, they increasingly emulate human-like reasoning with a level of independence that challenges traditional notions of human-AI interaction. The key characteristics of these agents include:
  • Planning: The ability to plan and execute multistep processes to achieve requested goals. For example, an agent can outline an essay, conduct online research, and draft the essay in a sequential manner (Ng 2023).
@@
These characteristics enable agents to complete sequential tasks with multiple iterations alongside human users, demonstrating chain-of-thought reasoning and providing explanations for their outputs. Continuous improvements in these abilities can eventually lead to the deployment of sophisticated agents that autonomously handle tasks alongside human users.
Agentic AI systems are fulfilling task specific applications such as for web searches, code execution, data processing, customer support, and other professional domains (Rahwan et al. 2019). Agentic AI’s ability to learn from all types of data and apply this knowledge to new and unforeseen scenarios make them suitable for a much wider range of applications in organizational and team processes. In this sense, Horton (2023) views these newly developed intelligent agents as implicit computational models of humans (namely, the so-called “homo silicus”). Their capability is not merely limited to resolving problems in a virtual environment (i.e., cyberspace) but also navigating the complexity and unpredictability of the physical world (Xu et al. 2024, Liu et al. 2024b).
In the following sections, we interchangeably use the terms “intelligent agents,” “agents,” and “AI” (in contrast to “traditional AI”) to refer to agents that possess the aforementioned intelligent characteristics.
@@ 2.2 Definition of Human–AI Teaming
According to Salas et al. (1995), a team is a group of two or more individuals who share common goals, role assignments, and interdependence. Typically, team dynamics (i.e., teaming) involve task completion within a specific context, application of specialized knowledge and skills, and performance under time, space, resource, cognitive, and other constraints. Teamwork includes two categories of behavior: task-work and team-work. Task-work involves actions performed by individual team members essential for carrying out their specific functions. Teamwork involves knowledge, skills, and attitudes that facilitate interaction among team members, ensuring coordination to achieve team goals (Morgan Jr et al. 1986). Essentially, teams operate as sociotechnical systems, where well-defined roles (akin to societal roles) and effective coordination mechanisms are fundamental to achieving high performance, especially in high-stakes environments and complex tasks (Salas 1992).
To initiate a team, it is necessary to confirm participant configurations, superordinate goals and priorities, the interdependence of teammate goals, sources of information and instruction for tasks, as well as factors such as team cohesion, communication, and coordination (Salas et al. 1995). Essentially, teamwork is a form of collaboration. Regardless of the levels of collaboration (Aaltonen et al. 2018), an essential characteristic of any team, including human-AI teams, is the presence of shared goals and the ability to complement each other in working on tasks (i.e., task assemblage) to achieve those goals (Rai et al. 2019).
@@
Agentic AI systems create new collaboration dynamics in teams. The infusion of such agents requires more cognitive input and coordination among human team members to ensure role fluidity and teamwork effectiveness (Gorman 2014). To facilitate teamwork between humans and intelligent agents, all participants need to adapt to changes in role specification withthe partnered intelligent agents(Zhang et al. 2021). Typically, human-AI teaming involves humans and intelligent agents interacting as team members in a context where the agent performs similar task-work and team-work functions as a human team member (McNeese et al. 2021). Based on the previous discussions, we present a formal definition of human-AI teaming when the focal emerging intelligent agents possess most of the characteristics defined in Section 2.1. In human-AI teaming, humans and intelligent agents collaborate to achieve shared goals. Human and AI team members dynamically coordinate their actions to effectively plan, facilitate, iterate, and evaluate. [^2] This follows Gao et al. (2024a) who define the interaction flow of human-LLM coworking.
@@
Human-AI teaming shares many similarities with human-human teaming, allowing for the extrapolation of existing interpersonal team management literature to this new context. For example, both forms of teaming involve shared objectives. Each member, whether human or AI, can be assigned specific tasks, contributing to both task-work and team-work. Teamwork involves multiple iterative interactions and feedback loops(Pei et al. 2024). Moreover, mutual learning is a key component and driver of team performance.
@@
There are, however, several key factors that distinguish AI agents from typical human teammates or traditional technical tools. While AI agents often operate under full control of humans regarding task assignments and (sub)objectives, AI’s advanced capabilities can flatten traditional team hierarchies (Tambe 2025, Wu et al. 2019a). Additionally, unlike human teammates, AI agents impose less social and emotional burden, as they do not impose the same level of interpersonal skills requirement as in human teamwork. This lack of emotional involvement can alter the dynamics within a team, reducing emotional support and potentially affecting cohesion (Xu et al. 2023). AI agents may remain more compliant in conflicts, lacking the nuanced emotional responses that humans might offer in tense situations. Another critical distinction lies in the issue of accountability. Unlike human teammates, AI cannot be entirely responsible for decisions or actions, which raises questions about how responsibility is distributed within a team. This lack of accountability can influence team dynamics and trust building. Moreover, AI’s potential for deceptive (possibly due to issues such as hallucination (Banerjee et al. 2024)) or overly accommodating behavior adds another layer of complexity, as it can sometimes cater to user expectations in misleading ways, raising concerns about the trustworthiness of AI in collaborative settings. These challenges call for new theoretical understanding of human-AI teaming factors that contribute to achieving shared objectives.
@@ 2.3 Theoretical Framework: An Extended Model of Team Situation Awareness
Team-related activities, as outlined in the Circumplex model of group tasks (Posner et al. 2005), typically involve generating ideas or plans, selecting the most suitable solution, negotiating a resolution to conflicts, and executing tasks. In the context of human-AI teaming, it becomes crucial to revisit and explore how team SA interacts with these complex tasks. AI’s capacity to rapidly process large volumes of data and generate insights can work for both role specification and fluidity to enhance team SA, potentially reshaping how teams handle decision-making, conflict resolution, and task execution. Additionally, as discussed earlier, the integration of AI introduces new complexities, such as the need for alignment between human and AI-generated awareness. As AI’s role in teams grows, revisiting the relationship between team SA and task performance is vital to ensure effective collaboration and to address potential gaps in mutual understanding.

### v2

@@
Artificial Intelligence (AI) is advancing at an unprecedented pace, with clear potential to enhance decision-making and productivity. Yet, the collaborative decision-making process between humans and AI remains underdeveloped, often falling short of its transformative possibilities. This paper explores the evolution of AI agents from passive tools to active collaborators in human-AI teams, emphasizing their ability to learn, adapt, and operate autonomously in complex environments. This paradigm shifts challenges traditional team dynamics, requiring new interaction protocols, delegation strategies, and responsibility distribution frameworks. Drawing on Team Situation Awareness (SA) theory, we identify two critical gaps in current human-AI teaming research: the difficulty of aligning AI agents with human values and objectives, and the underutilization of AI’s capabilities as genuine team members. Addressing these gaps, we propose a structured research outlook centered on four key aspects of human-AI teaming: formulation, coordination, maintenance, and training. Our framework highlights the importance of shared mental models, trust-building, conflict resolution, and skill adaptation for effective teaming. Furthermore, we discuss the unique challenges posed by varying team compositions, goals, and complexities. This paper provides a foundational agenda for future research and practical design of sustainable, high-performing human-AI teams.
@@
Artificial Intelligence; AI Agent; Human–AI Teaming; Situation Awareness; Research Outlook
@@ 1 Introduction
Artificial Intelligence (AI) is advancing at an unprecedented pace, transforming industries and reshaping daily life. From healthcare and finance to education and manufacturing, AI-based systems are increasingly supporting humans across various fields (Lu and Zhang 2025, Lou and Wu 2021, Berente et al. 2021). Yet, despite AI’s potential, collaborations between humans and AI remain ad hoc, often resulting in underwhelming performance (Teaming 2022). The promise of AI lies in its ability to integrate into teams to augment human capabilities, improve decision-making, enhance productivity, and drive innovation. Realizing the benefits, however, requires a more systematic examination of the novelties associated with human-AI teaming.
@@
In traditional human-human teaming, researchers have extensively studied shared awareness, communication, and collaboration, establishing frameworks around communication frequency, shared mental models, task complexity, and training. Similarly, human-system teaming has focused on the use of communication and collaboration tools, which support team coordination but remain distinct from the tasks themselves (DeSanctis and Gallupe 1987, Jerry Fjermestad 2000, Dennis and Wixom 2002). Implicit in all team-based interactions is the assumption that team members share common values and work together toward shared goals and objectives. The perspective that technology is a tool to enhance human efforts needs a refresh with the emergence of AI agents. Recent studies have shown that AI agents are evolving rapidly, with their capabilities doubling approximately every seven months since 2019 (Kwa et al. 2025). [^1] The concept of AI agents is not new. Russell and Norvig (2016) define them as entities that perceive their environment and act rationally. This paper builds on and modernizes that foundational view to reflect the advanced capabilities of today’s AI. Today’s AI agents have evolved into autonomous, iterative, and intelligent team members. They learn, reflect, and adapt within complex and dynamic environments. Moreover, they are no longer passive tools but active collaborators in human-AI teams in sustained interaction, emotional intelligence, and autonomous problem-solving. This marks a paradigm shift beyond the classical notion of rational action. In this regard, human-AI teaming introduces a new era of team science, requiring revised interaction protocols for AI agents, and new approaches to delegation, task execution, and responsibility sharing.
@@
In this paper, we examine the emerging landscape of human-AI teaming, where AI has moved beyond simple automation to become an intelligent, agentic artifact (Dennis et al. 2023, Gao et al. 2024b). The rapid rise of large language models (LLMs) such as Generative Pre-trained Transformer (GPT) has accelerated this transformation. AI-enabled agentic artifacts are no longer simply embedded into workflows; they are now intelligent agents capable of actively taking on mission-critical roles and tasks within teams (McKinsey 2023, 2024, Alavi 2024). These systems learn from interactions, adapt to changing contexts, and make independent decisions, transforming AI into an autonomous partner in collaborative efforts. The agentic shift highlights some gaps in the current information systems (IS) literature, which mainly treats AI as a tool rather than a nearly full-fledged teammate (Wang et al. 2019, Dennis et al. 2023, Lu and Zhang 2025).
@@
This paper aims to bridge these gaps by applying the well-established Team Situation Awareness (SA) theory to the context of human-AI teaming. We argue that integrating AI agents into teams not only requires designing systems that can work toward shared objectives but also requires a rethinking of how teams execute tasks alongside AI agents. By doing so, we develop an extended model of Team SA in human-AI teaming, and propose a structured outlook for future research. In particular, we propose a systematic framework for understanding human-AI teaming across four key dimensions: team formulation, coordination, social dynamics, and knowledge creation. First, effective team formulation requires shared mental models that align AI’s decision-making with human values and objectives, ensuring clarity in role specification while allowing flexibility for adaptation. Team coordination focuses on optimizing task delegation and human-AI interaction protocols, minimizing expectation misalignments and ensuring interpretable and reliable decisions. Team maintenance is crucial for trust-building, conflict resolution, and sustaining high-performance collaboration, particularly in managing AI’s limitations and accountability in decision-making. Finally, team training and evolution involve continuous learning mechanisms that enable AI to refine its capabilities while ensuring human members continue to create new knowledge. By structuring research around these four interconnected areas, we outline a roadmap for advancing human-AI teaming and addressing critical challenges such as trust, adaptability, accountability, and long-term sustainability in AI-integrated teams.
@@ 2.1 Emerging Agentic AI Systems
Information systems are beginning their transformation into agentic artifacts that “are no longer passive tools waiting to be used, are no longer always subordinate to the human agent, and can now assume responsibility for tasks with ambiguous requirements and for seeking optimal outcomes under uncertainty” (Baird and Maruping 2021). Modern AI systems, with their ability to “learn and act” independently of human instructions, are showing the promise to emulate human thought processes, behavior, and decision-making (Dennis et al. 2023, Durante et al. 2024). Although these systems may remain within the realm of Narrow AI, their capabilities now extend to many tasks that, until recently, required highly trained professionals.
Generative AI systems, such as ChatGPT, Midjourney, Claude, Llama, and Gemini, excel in tasks requiring nuanced judgment and adaptability — areas where classical AI systems often falter due to “brittleness” when encountering scenarios outside their predefined rules (McKinsey 2023, Dhar 2023). They now assist in domains previously dominated by human cognition and expertise (Zhou and Lee 2024, Susarla et al. 2023). For example, ChatGPT interprets ambiguous prompts to draft complex documents and hold meaningful conversations. Similarly, Midjourney generates original artwork from simple text, demonstrating creative problem-solving skills that go beyond rule-based automation. Importantly, these capabilities pave the way for the emergence of agentic AI systems—systems that not only assist but also act autonomously and collaborate dynamically with humans (Durante et al. 2024). Agentic AI engages in creative, coordinated tasks, takes initiative in decision-making, and adapts through iterative feedback. By generating context-aware outputs and exhibiting independent reasoning, they challenge traditional models of human-AI interaction.
The key characteristics of these agents include:
  • Planning: The ability to plan and execute multi-step processes to achieve requested goals. For example, an agent can outline an essay, conduct online research, and draft the essay in a sequential manner (Ng 2023).
@@
These characteristics enable agents to undertake sequential tasks with multiple iterations alongside human users, demonstrating chain-of-thought reasoning and providing explanations for their outputs. Continuous improvements in these abilities could eventually lead to the deployment of sophisticated agents that autonomously handle tasks alongside human users. Agentic AI systems are fulfilling domain-specific tasks such as web search, code execution, data processing, customer support, and other professional services (Rahwan et al. 2019). Their ability to learn from diverse data sources and generalize to new and unforeseen scenarios positions them for even broader applications in organizational and team-based processes. In this context, these intelligent agents have been conceptualized as implicit computational models of humans—referred to as “homo silicus” (Filippas et al. 2024). Notably, their capability is not merely limited to resolving problems in a virtual environment (i.e., cyberspace) but also navigating the complexity and unpredictability of the physical world (Xu et al. 2024, Liu et al. 2024b). In the following sections, we interchangeably use the terms “intelligent agents”, “agents”, and “AI” (in contrast to “traditional AI”) to refer to such systems that possess the aforementioned intelligent characteristics.
@@ 2.2 Definition of Human–AI Teaming
According to Salas (1992), a team is a group of two or more individuals who share common goals, role assignments, and interdependence. Typically, team dynamics (i.e., teaming) involve task completion within a specific context, application of specialized knowledge and skills, and performance under time, space, resource, cognitive, and other constraints. Teamwork includes two categories of behavior: task-work and team-work. Task-work involves actions performed by individual team members essential for carrying out their specific functions. Teamwork involves knowledge, skills, and attitudes that facilitate interaction among team members, ensuring coordination to achieve team goals (Morgan Jr et al. 1986). Essentially, teams operate as socio-technical systems, where well-defined roles (akin to societal roles) and effective coordination mechanisms are fundamental to achieving high performance, especially in high-stakes environments and complex tasks (Salas 1992).
To initiate a team, it is necessary to confirm participant configurations, superordinate goals and priorities, the interdependence of teammate goals, sources of information and instruction for tasks, as well as factors such as team cohesion, communication, and coordination (Salas 1992). At its core, teamwork is a form of collaboration. Regardless of the levels of collaboration (Aaltonen et al. 2018), an essential characteristic of any team, including human-AI teams, is the presence of shared goals and the ability to complement each other in working on tasks (i.e., task assemblage) to achieve those goals (Rai et al. 2019).
@@
Agentic AI systems create new collaboration dynamics in teams. The infusion of such agents requires more cognitive input and coordination among human team members to ensure role fluidity and teamwork effectiveness (Gorman 2014). To facilitate teamwork between humans and intelligent agents, all participants need to adapt to changes in role specification with the partnered agents (Zhang et al. 2021). Typically, human-AI teaming involves humans and intelligent agents interacting as team members in a context where the agent performs similar task-work and team-work functions as a human team member (McNeese et al. 2021). Based on the previous discussions, we present a formal definition of human-AI teaming when the focal intelligent agents possess most of the characteristics defined in Section 2.1. In human-AI teaming, humans and intelligent agents collaborate to achieve shared goals. Human and AI team members dynamically coordinate their actions to plan, facilitate, iterate, and evaluate. [^2] This follows Gao et al. (2024a) who define the interaction flow of human-LLM co-working.
@@
Human-AI teaming shares many similarities with human-human teaming, allowing for the extrapolation of existing interpersonal team management literature to this AI-related context. For example, both forms of teaming involve shared objectives. Each member, whether human or AI, can be assigned specific tasks, contributing to both task-work and team-work. Teamwork involves multiple iterative interactions and feedback loops (Pei et al. 2024). Moreover, mutual learning is a key component and driver of team performance.
@@
There are, however, several key factors that distinguish AI agents from typical human teammates or traditional technical tools. While AI agents can be subject to human control in terms of task assignments and (sub)objectives, their advanced capabilities can disrupt traditional team hierarchies by flattening authority structures and redistributing expertise (Tambe 2025, Wu et al. 2019a). In addition, unlike human teammates, AI agents impose less social and emotional burden, as they do not impose the same level of interpersonal skills requirement as in human teamwork. This lack of emotional involvement can alter the dynamics within a team, reducing emotional support and potentially affecting cohesion (Xu et al. 2023). AI agents may remain more compliant in conflicts, lacking the nuanced emotional responses that humans might offer in tense situations. A further distinction lies in the question of accountability. Unlike human teammates, AI cannot be entirely responsible for decisions or actions, which raises questions about how responsibility is distributed within a team. This ambiguity can complicate trust-building and decision-making processes within a team. Moreover, AI’s potential for deceptive (possibly due to issues such as hallucination (Banerjee et al. 2024)) or overly accommodating behavior adds another layer of complexity, as it can sometimes cater to user expectations in misleading ways, creating false confidence and undermining the reliability of collaboration. These challenges call for a new theoretical understanding of human-AI teaming factors that contribute to achieving shared objectives.
@@ 2.3 Theoretical Framework: An Extended Model of Team Situation Awareness
Team-related activities, as outlined in the Circumplex model of group tasks (Posner et al. 2005), typically involve generating ideas or plans, selecting the most suitable solution, negotiating a resolution to conflicts, and executing tasks. In the context of human-AI teaming, it becomes crucial to revisit and explore how team SA interacts with these complex tasks. AI’s capacity to rapidly process large volumes of data and generate insights can work for both role specification and fluidity to enhance team SA, potentially reshaping how teams handle decision-making, conflict resolution, and task execution. In addition, as discussed earlier, the integration of AI introduces new complexities, such as the need for alignment between human and AI-generated awareness. As AI’s role in teams grows, revisiting the relationship between team SA and task performance is vital to ensure effective collaboration and to address potential gaps in mutual understanding.